OpenStudy (anonymous):

A fair coin is tossed 3 times in a row. What is the probability that heads appears on only the last toss?

OpenStudy (anonymous):

@Kitten_is_back

OpenStudy (anonymous):

@YoungStudier

OpenStudy (anonymous):

#medal given

OpenStudy (kitten_is_back):

okay

OpenStudy (anonymous):

soooo..... what is the answer?

OpenStudy (kitten_is_back):

holf on

OpenStudy (anonymous):

@ajo21

OpenStudy (anonymous):

@Error1603

OpenStudy (kitten_is_back):

1/8 is the answer

OpenStudy (error1603):

Your right but how did you get it?

OpenStudy (error1603):

P(2 heads) = 3/8 = 1/4 = 0.375 = 37.5% (HHT, HTH, THH) P(at least 2 heads) = P(2 heads) + P(3 heads) = 3/8 + 1/8 = 4/8 = 0.5 = 50% P(at least 1 head) = P(1H) + P(2H) + P(3H) = 3/8 + 3/8 + 1/8 = 7/8 = 0.875 = 87.5% P(1 head) = 3/8 = 0.375 = 37.5% (HTT, THT, TTH) P(1 head on last toss) = 1/8 = 0.125 = 12.5% (TTH) @klc097

OpenStudy (kitten_is_back):

just poped up in m u hwad